OpenStudy (anonymous):

Which of the following properties is not used to simplify the expression in the steps below? y plus 5 times x minus 5 times the quantity x minus y y plus 5 times x minus 5 times x plus 5 times y y plus zero times x plus 5 times y y plus zero plus 5 times y zero plus y plus 5 times y zero plus 1 times y plus 5 times y zero plus 6 times y 6 times y Commutative Property of Addition Zero Property of Multiplication Identity Property of Addition Associative Property of Multiplication

myininaya (myininaya):

y+5x-5(x-y) = y+5x-5x+5y = y+0x+5y = y+0+5y = 0+y+5y = 0+1y+5y = 0+6y = 6y First of all it helps to write using math symbols! Is it pretty like this on your paper or it actually written out in words?

OpenStudy (anonymous):

SO IS IT Commutative Property of Addition, Zero Property of Multiplication, Identity Property of Addition ,Associative Property of Multiplication

myininaya (myininaya):

y+5x-5(x-y) = y+5x-5x+5y <---Distributive property was used! = y+0x+5y <---Which is this one that says 5x-5x=0? = y+0+5y <---And then this one 0*x=0? = 0+y+5y <---They changed order a+b=b+a, which is this? = 0+1y+5y <--- y=1y another property was used. = 0+6y <---Combined like terms = 6y <---One more property was used 0+6y=6y, which was it?

myininaya (myininaya):

a+b=b+a which property is this?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\psi\]

myininaya (myininaya):

Ok a+b=b+a is the commutative property of addition

myininaya (myininaya):

The order changes and its addition then the commutative property was used

myininaya (myininaya):

Do you know what the associative property of addition is?